.Calculate the rate of heat energy emitted by 65 m2 of a polished iron surface with an emissivity of 0.06 and a surface temperature of 40*C


1
Expert's answer
2021-05-11T09:07:56-0400

According to Stephan-Boltzman law:


"\frac Qt=\sigma eA(T_2^4-T_1^4),"


where

"\sigma" - Stephan-Boltzmann constant,

e - emissivity,

A - area, m2

T2 - surface temperature, K,

T1 - temperature of the environment, 300 K.


"\frac Qt=5.67·10^{-8}·0.06·65(313^4-300^4)=331\text{ W}."